# Anti-correlation between excitations and locally-favored structures in glass-forming systems

### v1

@@ Abstract
Dynamics that are microscopic in space and time, so-called excitations, are considered the elementary unit of relaxation in the Dynamic Facilitation (DF) theory of the glass transition. Meanwhile, geometric motifs known as locally favored structures (LFS) are associated with vitrification in many glassformers. Recent work indicates that the probability of particles both found in locally favored structures (LFS) and excitations decreases significantly upon supercooling suggesting that there is an anti-correlation between them. [Ortlieb et al, Nature Commun. 14, 2621 (2023)]. However, the spatial relationship between excitations and LFS remains unclear. By employing state-of-the-art GPU computer simulations and colloid experiments, we analyze this relationship between LFS and excitations in model glassformers. We demonstrate that there is a spatial separation between the two in deeply supercooled liquids. This may be due to the fact that LFS are well-packed, thus they are solid and stable.
@@ I Introduction
The process of vitrification, whereby a liquid solidifies without crystallizing, remains a major challenge in condensed matter. There are a variety of theories postulated, which provide equally good descriptions of the observed dynamic slowdown of some fourteen orders of magnitude in relaxation time with respect to the normal iquid Berthier and Biroli (2011). Some are even understood to be incompatible, for example those which relate the dynamic slowdown to a thermodynamic transition to a putative amorphous state with sub–extensive configurational entropy known as the ideal glass Adam and Gibbs (1965); Lubchenko and Wolynes (2007) while others suppose that the glass transition is a dynamical phenomenon Chandler and Garrahan (2010); Speck (2019). Still others relate the glass transition to the emergence of geometric motifs associated with local order. These so-called locally favored structures (LFS) may be amorphous Tarjus et al. (2005); Royall and Williams (2015) or crystalline Leocmach and Tanaka (2012).
Particle–resolved studies, using computer simulation or optical imaging of colloids Hunter and Weeks (2012); Ivlev, A. and Lowen, H and Morfill, G. E. and Royall, C. P. (2012); Royall et al. (2023); Gokhale et al. (2016) promise the level of data to discriminate between the predictions made by these competing theoretical descriptions. Yet until recently, these are limited to the weakly supercooled regime of around four orders of magnitude increase in relaxation time. Now this regime is well–described by the Mode–Coupling theory Charbonneau and Reichman (2005). Particle–resolved data obtained at state points which are more deeply supercooled than the crossover of Mode-Coupling Theory (MCT) is needed to make progress.
@@
Recently, such data has become available, using SWAP Monte Carlo Ninarello et al. (2017); Scalliet et al. (2022), and GPU processors Bailey et al. (2015); Ortlieb et al. (2023). Colloid experiments too have passed the MCT crossover, using smaller particles Brambilla et al. (2009); Hallett et al. (2018); Ortlieb et al. (2023). However, rather than providing clear support for one particular theoretical approach, in fact these new data seem to support multiple theories. One of these is dynamic facilitation theory Chandler and Garrahan (2010); Speck (2019); Hasyim and Mandadapu (2024) which predicts that relaxation occurs by excitations which are microscopic in space and time. These are indeed found at the population, size and duration predicted in simulations Keys et al. (2011); Ortlieb et al. (2023); Hasyim and Mandadapu (2021, 2024) and experiments Gokhale et al. (2014, 2016). On the other hand the co-operatively re-arranging regions (CRRs) predicted by the thermodynamics-based approaches of the Random First–Order Transition theory (RFOT) Lubchenko and Wolynes (2007) and Adam-Gibbs theory Adam and Gibbs (1965) are predicted to grow in size and massively in timescale upon supercooling. These too are found, consistent with theory and in particular the prediction of their compaction at deep supercooling is upheld Nagamanasa et al. (2015); Ortlieb et al. (2023). In addition, the predicted drop in configurational entropy has been robustly found in a number of studies Turci et al. (2017); Berthier et al. (2017); Hallett et al. (2018); Berthier et al. (2019). It has even been suggested that these two approaches may be reconciled Royall et al. (2020); Ortlieb et al. (2023).
One piece of the jigsaw which has received rather little attention is the relationship between locally favoured structures and dynamic facilitation theory Chandler and Garrahan (2010); Speck (2019). LFS play an important role in frustration based theories Tarjus et al. (2005), and a key piece of evidence that there is at least some structural component to dynamic facilitation comes from the observation that the time–averaged population of LFS can drive a dynamical phase transition which is a key ingredient of the theory Chandler and Garrahan (2010); Speck et al. (2012); Turci et al. (2017). Very recently, a strong anti–correlation between LFS and particles in excitations has been found Ortlieb et al. (2023). This contrasts with previous work which could only access weaker supercooling and did not significantly pass the mode-coupling crossover Malins et al. (2013a); Royall and Williams (2015). Amongst earlier work, some studies claimed a significant relationship between local order and dynamic heterogeniety Kawasaki et al. (2007); Leocmach and Tanaka (2012); Tamborini et al. (2015), while other work found rather little Charbonneau et al. (2012). While it is likely that the means of determining the structure may be important Richard et al. (2020); Tong and Tanaka (2018); Cubuk et al. (2015); Malins et al. (2013b), in any case, structure-dynamics relationships in the weakly super–cooled regime in 3d have been found to be model–dependent Hocky et al. (2014).
@@
The fact that it is now possible to study glass-formers at the single-particle level in real space at deeper supercooling than the Mode-Coupling crossover opens the way to probe the role of LFS in this newly accessible dynamical regime. The anti-correlation between LFS and excitations Ortlieb et al. (2023) suggests that it may be interesting to investigate further the relationship between LFS and excitations of dynamic facilitation, particularly at deeper supercooling, which forms the topic of this work. Moreover, if one imagines that the LFS are somehow stable relative to other parts of the system, then upon supercooling they might be expected to “expel” excitations. Here, we investigate this hypothesis using GPU computer simulations of a model glassformer and colloid experiments.
@@
This paper is organised as follows. In the following methods (Sec. II), we provide some details of our computer simulations, experiments and analysis. We present our results in Sec. III. Our findings are discussed in Sec. IV and we conclude in Sec. V. Supporting information is provided in Sec. VI.
@@ II Methods
Figure 1: Characteristics of the systems considered. (a) Angell plot of relaxation time $\tau_{\alpha}$ with temperature $T$ for Kob-Andersen binary (KA) mixture and compressibility $Z$ for the hard-sphere colloidal system, scaled with Vogel-Fulcher-Tamman (VFT) parameters. Data for KA 2:1 and 3:1 are from Ortlieb et al. (2023) and Exp. are from Royall et al. (2018). The black line is the fit to the Vogel Fulcher-Tamman equation (Eq. 3) for KA 2:1. The fitting parameters are shown in the Supporting Information (SI), Sec. VI, Table S1. Blue shading indicates supercooling beyond the mode-coupling crossover. Data points circled in black are used in the following analysis. (b) The proportion of long-lived LFS. Upper left: the geometric form of a bicapped square antiprism, the LFS for the KA mixture. Lower right: the LFS for the experimental colloidal system, the defective icosahedron. (c) The concentration of excitations. The fitting parameters are shown in the SI, Table S2.
@@ II.1 Details of computer simulations
We study the Kob-Andersen (KA) binary Lennard-Jones mixture at 2:1 and 3:1 compositions with density $\rho=1.4$. The system size is chosen to be N=10002 and 10000 for 2:1 and 3:1 respectively. The interactions in the KA mixture are defined by the Lennard-Jones potential
@@
with parameters $\sigma_{AA}=1,\sigma_{AB}=0.80,\sigma_{BB}=0.88$, and $\epsilon_{AA}=1,\epsilon_{AB}=1.50,\epsilon_{BB}=0.50(\alpha,\beta=A,B)$, and we employ $m_{A}=m_{B}=1$ (we refer to these as LJ units). For both compositions, the density was fixed to 1.4.
Computer simulations were carried out using Roskilde University Molecular Dynamics (RUMD). This is a Molecular Dynamics code which takes advantage of multiple GPU cores to achieve high performance Bailey et al. (2017). Thus, RUMD makes it possible to probe the system at very low temperatures (e.g. $T=0.48$, where the structural relaxation time $\tau_{\alpha}>10^{6}$). Please see Ortlieb et al. (2023) for further details of the RUMD simulations. After the system reaches equilibrium, we use the LAMMPS package Plimpton (1995); Thompson et al. (2022) to produce shorter trajectories of 1000 LJ time units which we sample with a time interval of 1 LJ time unit in order to identify excitations see below, Sec. II.3.
To obtain the inherent structures (IS), we split the obtained trajectory into frames and apply the LAMMPS [minimize] command to each frame. We produce 10 independent IS trajectories for each temperature [0.48,0.49,0.50,0.52,0.55], with the timescale of 1000 LJ time units for further analysis. All analysis here is carried out using the inherent state, as we found this provided a stronger signal.
@@ II.2 Experimental details
We carry out confocal microscopy experiments with colloidal particles which we track at the single particle level in space and time. These particles closely approximate the hard sphere model Royall et al. (2023, 2018). We use fluorescently labeled density and refractive index matched colloids of sterically stabilized polymethyl methacrylate. The diameter of the colloids was $\sigma^{\mathrm{exp}}$= 3.23 $\mu$m and the polydispersity was 6% which is sufficient to suppress crystallization here. The Brownian time to diffuse a radius here is $\tau_{B}=6.06$s. The particles were labelled with the fluorescent dye 3,3’-dioctadecyloxacarbocyanine perchlo- rate (DiOC18). Further details are available in Royall et al. (2018). In this system, the locally favoured structure is the defective icosahedron Royall et al. (2015, 2018). This structure is pictured in Fig. 1(b), and (unlike the antiprism of the KA model), it does not contain a full shell around a central particle. Below, we shall consider the distance from a given particle to the center of the LFS. For the defective icosahedron we take a particle close to the centre of mass around $0.45\sigma$, as illustrated in Fig. 1(b).
In the case of hard spheres, the glass transition is obtained via compression, or increasing the volume fraction $\phi$. The reduced pressure $Z=\beta P/\rho$ where $\beta=1/k_{B}T$ the thermal energy, $P$ is pressure and $\rho$ is number density is a convenient quantity with which to express the state point Berthier and Witten (2009). Here $Z$ is determined from the Carnahan–Starling relation.
$Z=\frac{1+\phi+\phi^{2}-\phi}{(1-\phi)^{3}}$ (2)
where $\phi$ is the particle volume fraction. Although the colloidal system approaches its glass transition via compression rather than cooling, to facilitate comparison with the simulations (and, more generally, molecular systems), we nevertheless take the liberty of referring to the colloidal system as being supercooled Royall et al. (2023).
@@ II.3 Detecting excitations
In the case of the computer simulations, we use the algorithm proposed by Ortlieb. et al  Ortlieb et al. (2023) to detect excitations, although here we apply the algorithm to inherent state configurations. For the experiments we make a few small changes as described in the procedure below. The algorithm is executed as follows:
@@
    Dividing the trajectory. We divide our 1000 LJ time unit trajectory into 5 sub-trajectories of length $t_{a}^{\mathrm{sim}}=200$ in each piece in the case of the simulations. For the experiments, we take $t_{a}^{\mathrm{exp}}=200\tau_{B}$.
@@
    Testing whether the particle has committed to a new position. The average particle positions of the first and the last $t_{a}$ of the sub-trajectory are compared. If the difference between these positions is smaller than the predefined threshold $a$, the particle is rejected. Here we choose $a=0.5\sigma$.
@@
    Determining the time at which the excitation takes place. This time is denoted as $t_{0}$, and the corresponding duration $\Delta t$ and displacement $\Delta x$ are found using a hyperbolic tangent fit. We set a sliding window to apply the hyperbolic tangent fit, by initially considering every frame as the center of the fit. The central position of the local optimal fit is considered to be where excitations occur ($t_{0}$). It should be noted that it turns out that there is a tiny chance ($<5\%$) of one particle exhibiting more than once excitation along the sub-trajectory. So, if the central position $t_{0}$ of two local optimal fit is larger than $t_{a}$, we keep both.
@@
    Excluding slow and steady movement. We exclude particles with duration $\Delta t>\frac{3}{4}t_{a}$ and displacement $\Delta x<a$.
@@
    Excluding the return of particles to the original position. In the case of the simulations, we use the same simulation method, we run the system for another 1000 LJ time units and check the final displacement of those particles (to be precise, the comparison of the average position of the first and last $t_{a}$ of the trajectory of 2000 LJ time units). We exclude particles with displacement smaller than $a$. In the case of the experiments, we consider the remainder of the trajectory.
@@ II.4 Identifying long-lived LFS
We use the Topological Cluster Classification (TCC) algorithm to probe LFS in the Kob Andersen mixtures and hard sphere colloids Malins et al. (2013b). The LFS has been previously identified for the Kob Andersen model as the bicapped square antiprism and defective icosahedron for hard spheres  Coslovich and Pastore (2007); Malins et al. (2013a); Crowther et al. (2015); Dunleavy et al. (2015); Royall et al. (2015). These LFS are comprised of 11 and 10 particles respectively and are indicated in Fig. 1(b). The initial step of the TCC algorithm involves identifying bonds between neighboring particles. These bonds are detected using a modified Voronoi method, where a maximum bond length cutoff of $r_{c}=2.0$ is applied for all types of interactions (AA, AB, and BB). Additionally, for KA the parameter $f_{c}$ is set to unity to control the identification of four-membered rings as opposed to three-membered rings, while $f_{c}=0.82$ for hard spheres and all particles are treated equally, ie polydispersity is neglected following previous work Malins et al. (2013a); Dunleavy et al. (2015); Royall et al. (2015).
After identifying the LFS, we select those which survive over 80% of the duration of the sub-trajectory (i.e. longer than 800 LJ time units). We choose the 80% criterion on the basis of the probability distribution of LFS lifetime, which serves as a turnover point [Fig. S2]. We consider those clusters to be long-lived LFS. It should be noted that here we do not require these clusters to be in LFS for 800 LJ time units continuously.
In the following, we consider exclusively long-lived LFS following the discussion above.
Figure 2: Anti-correlation between LFS and excitations. Thermalised/instantaneous data from Ortlieb et al. (2023) (red dashed line). Solid lines correspond to the anti-correlation between long-lived LFS and excitations in the inherent states.

### v2

@@ Abstract
Dynamics that are microscopic in space and time, in which particles commit to a position, so-called excitations, are considered the elementary unit of relaxation in the Dynamic Facilitation (DF) theory of the glass transition. Meanwhile, geometric motifs known as locally favored structures (LFS) are associated with vitrification in many glassformers. Recent work indicates that the probability of particles found both in locally favored structures (LFS) and excitations decreases significantly upon supercooling suggesting that there is an anti-correlation between them [Ortlieb et al, Nature Commun. 14, 2621 (2023)]. However, the spatial relationship between excitations and LFS remains unclear. By employing state-of-the-art GPU computer simulations and colloid experiments, we analyze this relationship between LFS and excitations in model glassformers. We demonstrate that there is a spatial separation between the two in deeply supercooled liquids. This may be due to the fact that LFS are well-packed, thus they are relatively stable.
@@ I Introduction
The process of vitrification, whereby a liquid solidifies without crystallizing, remains a major challenge in condensed matter. There are a variety of theories postulated, which provide equally good descriptions of the observed dynamic slowdown of some fourteen orders of magnitude in relaxation time with respect to the normal liquid Berthier and Biroli (2011). Some are even understood to be incompatible, for example those which relate the dynamic slowdown to a thermodynamic transition to a putative amorphous state with sub–extensive configurational entropy known as the ideal glass Adam and Gibbs (1965); Lubchenko and Wolynes (2007) while others suppose that the glass transition is a dynamical phenomenon Chandler and Garrahan (2010); Speck (2019). Still others relate the glass transition to the emergence of geometric motifs associated with local order. These so-called locally favored structures (LFS) may be amorphous Tarjus et al. (2005); Royall and Williams (2015) or crystalline Leocmach and Tanaka (2012).
@@
Particle–resolved studies, using computer simulation or optical imaging of colloids Hunter and Weeks (2012); Ivlev, A. and Lowen, H and Morfill, G. E. and Royall, C. P. (2012); Royall et al. (2023); Gokhale et al. (2016) promise the level of data to discriminate between the predictions made by these competing theoretical descriptions. Yet until recently, these are limited to the weakly supercooled regime of around four orders of magnitude increase in relaxation time. Now this regime is well–described by the Mode–Coupling theory Reichman and Charbonneau (2005). Particle–resolved data obtained at state points which are more deeply supercooled than the crossover of Mode-Coupling Theory (MCT) is needed to make progress. Recently, such data has become available, using SWAP Monte Carlo Ninarello et al. (2017); Scalliet et al. (2022), and GPU processors Bailey et al. (2017); Ortlieb et al. (2023). Colloid experiments too have passed the MCT crossover, using smaller particles Brambilla et al. (2009); Hallett et al. (2018); Ortlieb et al. (2023). However, rather than providing clear support for one particular theoretical approach, in fact these new data seem to support multiple theories. One of these is dynamic facilitation theory Chandler and Garrahan (2010); Speck (2019); Hasyim and Mandadapu (2024) which predicts that relaxation occurs by excitations which are microscopic in space and time. These are indeed found at the population, size and duration predicted in simulations Keys et al. (2011); Ortlieb et al. (2023); Hasyim and Mandadapu (2021, 2024) and experiments Gokhale et al. (2014, 2016). On the other hand the cooperatively re-arranging regions (CRRs) predicted by the thermodynamics-based approaches of the Random First–Order Transition theory (RFOT) Lubchenko and Wolynes (2007) and Adam-Gibbs theory Adam and Gibbs (1965) are predicted to grow in size and massively in timescale upon supercooling. These too are found, consistent with theory and in particular the prediction of their compaction at deep supercooling is upheld Hima Nagamanasa et al. (2015); Ortlieb et al. (2023). In addition, the predicted drop in configurational entropy has been robustly found in a number of studies Turci et al. (2017); Berthier et al. (2017); Hallett et al. (2018); Berthier et al. (2019). It has even been suggested that these two approaches may be reconciled Royall et al. (2020); Ortlieb et al. (2023).
[…]
@@
The fact that it is now possible to study glass-formers at the single-particle level in real space at deeper supercooling than the Mode-Coupling crossover opens the way to probe the role of LFS in this newly accessible dynamical regime. The anti-correlation between LFS and excitations Ortlieb et al. (2023) motivates us to further investigate the relationship between LFS and excitations of dynamic facilitation, particularly at deeper supercooling. This investigation forms the topic of this work. Moreover, if one imagines that the LFS are somehow stable relative to other parts of the system, then upon supercooling they might be expected to “expel” excitations. Here, we investigate this hypothesis using GPU computer simulations of a model glassformer and colloid experiments.
@@
This paper is organised as follows. In Sec. II we describe our computational and experimental models and methods. We present our results in Sec. III. Our findings are discussed in Sec. IV and we conclude in Sec. V. Supporting information is provided in Sec. VI.
@@ II Methods
Figure 1: Characteristics of the systems considered. (a) Relaxation time $\tau_{\alpha}$ with inverse temperature $T$ for Kob-Andersen binary (KA) mixtures (data from Ortlieb et al. (2023)) and compressibility $Z$ for the hard-sphere colloidal system (data from Royall et al. (2018)), scaled with Vogel-Fulcher-Tamman (VFT) parameters given in Table 1. The black line is the VFT fit for KA 2:1. We perform our analysis on the data points circled in black, below the mode-coupling crossover (blue shading). (b) Population of long-lived LFS: the bicapped square antiprism for KA mixtures (top left), and the defective icosahedron for the experimental colloidal system (bottom right). The pink particle is considered the central particle, while the black point is the geometric centre of the LFS. (c) Scaling of the excitation concentration $c_{a}$, with the fitting parameters given in Table 2.
@@ II.1 Details of computer simulations
We study the Kob-Andersen (KA) binary Lennard-Jones mixture at 2:1 and 3:1 compositions with particle density $\rho=1.4$. The system size is $N$=10002 for the 2:1 composition and $N$=10000 for 3:1. The interactions in the KA mixture are defined by the Lennard-Jones potential
@@
with parameters $\sigma_{AA}=1$, $\sigma_{AB}=0.80$, $\sigma_{BB}=0.88$, and $\epsilon_{AA}=1$, $\epsilon_{AB}=1.50$, $\epsilon_{BB}=0.50$ ($\alpha,\beta=A,B)$. We employ $m_{A}=m_{B}=1$. Times are expressed in units of $\sigma_{AA}\sqrt{m_{A}/\epsilon_{AA}}$.
Computer simulations were carried out using Roskilde University Molecular Dynamics (RUMD). This is a molecular dynamics code which takes advantage of multiple GPU cores to achieve high performance Bailey et al. (2017). RUMD makes it possible to probe the system at very low temperatures in equilibrium conditions (e.g. $T=0.48$, where the structural relaxation time $\tau_{\alpha}>10^{6}$). More details on the RUMD simulations are provided in  Ortlieb et al. (2023).
We simulate the systems in equilibrium conditions at temperature $T=$ 0.48, 0.49, 0.50, 0.52, 0.55 for the 2:1 composition, and $T=$ 0.68, 0.7, 0.73, 0.75, 0.8 for the 3:1. Once thermalization is reached at these temperatures, we use the LAMMPS package Plimpton (1995); Thompson et al. (2022) to produce short trajectories of 1000 LJ time units. We sample configurations at a time interval of 1 LJ time unit along the trajectory in order to identify excitations and LFS. For each temperature, we produce 10 independent trajectories, over which we average results.
We employ the conjugate gradient method to find the inherent structure (IS) of each sampled configuration. Our analysis is carried out using the IS trajectories, as we found this choice provides a stronger signal than the thermalized configurations.
@@ II.2 Experimental details
We carry out confocal microscopy experiments with colloidal particles which we track at the single particle level in space and time. These particles closely approximate the hard sphere model Royall et al. (2023, 2018). We use fluorescently labeled density and refractive index matched colloids of sterically stabilized polymethyl methacrylate. The diameter of the colloids is $\sigma^{\mathrm{exp}}$= 3.23 $\mu$m and the polydispersity is 6% which is sufficient to suppress crystallization. The Brownian time to diffuse a radius is $\tau_{B}=6.06$ s. The particles were labelled with the fluorescent dye 3,3’-dioctadecyloxacarbocyanine perchlo- rate (DiOC18). Further details are available in Royall et al. (2018). In this system, the locally favoured structure is the defective icosahedron Royall et al. (2015, 2018). This structure is pictured in Fig. 1(b).
The glass transition of hard spheres is obtained via compression, or increasing the volume fraction $\phi$. A convenient quantity to express the state point is the reduced pressure $Z=\beta P/\rho$ Berthier and Witten (2009), with $\beta=1/k_{B}T$ the inverse thermal energy, $P$ the pressure and $\rho$ the number density. Here $Z$ is determined from the Carnahan–Starling relation
$Z=\frac{1+\phi+\phi^{2}-\phi}{(1-\phi)^{3}}~{}~{}.$ (2)
Although the colloidal system approaches its glass transition via compression rather than cooling, to facilitate comparison with the simulations and, more generally, molecular systems, we still refer to the colloidal system as being supercooled Royall et al. (2023).
@@ II.3 Detecting excitations
We first describe the procedure to detect excitations in simulations and experiments, based on the algorithm proposed by Ortlieb et al.  Ortlieb et al. (2023). We analyse molecular dynamics data applying this algorithm to the inherent state trajectory, instead of the thermal one. To analyze experimental trajectories, we make a few small changes as described in the procedure below. The algorithm to detect excitations proceeds as follows:
@@
    Dividing the trajectory. We divide each 1000 LJ time unit trajectory into 5 sub-trajectories of length $t_{a}$. We use $t_{a}=200$ LJ time units for simulations, and $t_{a}=200\tau_{B}$ for experiments.
@@
    Testing whether the particle has committed to a new position. For each particle, we compare its average position in the first and last sub-trajectories. If the difference between these positions is smaller than a predefined threshold $a$, the particle is rejected, i.e. not considered an excitation. We choose $a=0.5\sigma$.
@@
    Determining the time at which the excitation takes place. This time is denoted as $t_{0}$, the excitation duration $\Delta t$ and its displacement $\Delta x$. They are found using a hyperbolic tangent fit to the particle position as a function of time. We set a sliding window on which the fit is applied, initially considering every frame as the center of the time window (corresponding to a potential jump in particle position, equivalently the excitation location). The excitation time $t_{0}$ corresponds to the center of the time window for which the fit is locally optimal. We find that there is a small probability ($<5\%$) that one particle exhibits multiple excitations. Thus, if two locally optimal fits yield excitation times $t_{0}$ separated by more than $t_{a}$, we keep both as excitations.
@@
    Excluding slow and steady movement. We do not consider as excitations particles that yield a duration $\Delta t>\frac{3}{4}t_{a}$ and displacement $\Delta x<a$.
@@
    Excluding the return of particles to the original position. For the simulations, we continue the MD trajectory for another 1000 LJ time units and check the final displacement of those particles. More precisely, we compare their average position in the first and last $t_{a}$ of the total 2000 LJ time units trajectory. We exclude particles with a displacement smaller than $a$. For the experiments, we consider the remainder of the trajectory.
@@
In addition to identifying individual excitations along trajectories, we also compute the concentration of excitations $c_{a}$ defined as the number of excitation found per trajectory (containing 1000 configurations).
@@ II.4 Identifying long-lived LFS
We now turn to the identification of LFS. We use the Topological Cluster Classification (TCC) algorithm to probe LFS in the Kob-Andersen mixtures and hard sphere colloids Malins et al. (2013b). The LFS has been previously identified for the Kob-Andersen model as the bicapped square antiprism and defective icosahedron for hard spheres  Coslovich and Pastore (2007); Malins et al. (2013a); Crowther et al. (2015); Dunleavy et al. (2015); Royall et al. (2015). These LFS are comprised of 11 and 10 particles respectively and are illustrated in Fig. 1(b). The initial step of the TCC algorithm involves identifying bonds between neighboring particles. These bonds are detected using a modified Voronoi method, where a maximum bond length cutoff of $r_{c}=2.0$ is applied for all types of interactions (AA, AB, and BB). Additionally, for KA the parameter $f_{c}$ is set to unity to control the identification of four-membered rings as opposed to three-membered rings, while $f_{c}=0.82$ for hard spheres and all particles are treated equally (we neglect polydispersity, following previous work Malins et al. (2013a); Dunleavy et al. (2015); Royall et al. (2015)).
After identifying the LFS, we are interested in selecting only those that are long-lived. In simulations, we define as long-lived the LFS that survive over 80% of the duration of the trajectory, i.e. longer than 800 LJ time units, not necessarily consecutively. We choose the 80% criterion on the basis of the probability distribution of LFS lifetime, which serves as a turnover point, see Fig. S1(a,b). For experiments, we consider clusters that survive over 40% time of the whole trajectory to be long-lived LFS, according to the LFS lifetime distribution Fig. S1(c).
In the following, our analysis is based on long-lived LFS.
Figure 2: Anti-correlation between LFS and excitations. Thermalized/instantaneous data from Ortlieb et al. (2023) (black dashed line with red data points). Experimental data with long-lived LFS is shown with the blue data points. Solid lines correspond to the anti-correlation between long-lived LFS and excitations in the inherent states.